i have to solve the problems using system of equations..
1.x+y=6 2. 2x-3y=-13
y=2x y=2x+7

Answer by mathick(4)

Let's start with #1:
x + y = 6
y = 2x.



There are a few ways to solve these, but the substitution method is often easiest.



1. Substitute 2x in for y in the 1st equation: x + (2x) = 6.
(This can be done because the 2nd equation says y and 2x are interchangeable).



2. Simplify: 3x = 6.



3. Solve for x: x = 2.



4. Plug 2 in for x in the 1st equation: 2 + y = 6.



5. Solve for y: y = 4.



6. Final Answer: (2, 4). (x = 2, y = 4).



That's the substitution method - the next example has a few more steps:



2x-3y=-13
y=2x+7



1. Substitute 2x + 7 in for y in the 1st equation: 2x - 3(2x + 7) = -13.
(This can be done because the 2nd equation says y and 2x + 7 are interchangeable).



2. Simplify: -4x - 21 = -13.



3. Solve for x: x = -2.



4. Plug -2 in for x in the 1st equation: 2(-2) - 3y = -13.



5. Solve for y: y = 3.



6. Final Answer: (-2, 3). (x = -2, y = 3).